OpenStudy (anonymous):

If sin0=3/5, then tan0=___.How do I do this?

OpenStudy (lgbasallote):

first think of sin 0 and tan 0 as sin x and tan x..okay?

OpenStudy (amistre64):

2 ways that i can think of off hand

OpenStudy (amistre64):

draw a representative right triangle using the definitions of sin and tan to solve the sides or using inverse trig functions and a calculator

OpenStudy (lgbasallote):

i was thinking hypotenuse adjacent and opposite..

OpenStudy (lgbasallote):

seems i was wrong..

OpenStudy (anonymous):

ok so its: \[\sin(x)=3/5 \tan(x)=___ \]